OpenStudy (anonymous):

What is the factorization of the expression below? 4x2 - 81

OpenStudy (jazzyluvsu123):

using the difference of squares: (a2−b2)=(a+b)(a−b) Think of 4x2=2x×2x and 81=9×9 If 4x^2=a^2 and 81=b^2, you get a=2x and b=9 So, factored would be (2x+9)(2x-9)

OpenStudy (jazzyluvsu123):

Hope this helps ;P